# Probing quantum chaos through singular-value correlations in sparse non-Hermitian SYK model

### v1

@@ Abstract
Utilizing singular value decomposition, our investigation focuses on the spectrum of the singular values within a sparse non-Hermitian Sachdev-Ye-Kitaev (SYK) model. Unlike the complex eigenvalues typical of non-Hermitian systems, singular values are inherently real and positive. Our findings reveal a congruence between the statistics of singular values and those of the analogous Hermitian Gaussian ensembles. An increase in sparsity results in the non-Hermitian SYK model deviating from its chaotic behavior, a phenomenon precisely captured by the singular value ratios. Our analysis of the singular form factor ($\upsigma$FF), analogous to the spectral form factor (SFF) indicates the disappearance of the linear ramp with increased sparsity. Additionally, we define singular complexity, inspired by the spectral complexity in Hermitian systems, whose saturation provides a critical threshold of sparseness. Such disintegration is likely associated with the breakdown of the existing holographic dual for non-Hermitian systems.
@@
Introduction: Spectral statistics serve as a critical tool for probing the energy levels in quantum systems, offering insights into the dynamics within quantum chaos and Random Matrix Theory (RMT) Mehta (1991); Bohigas et al. (1984); Wigner (1955, 1957); Dyson (1962a, b); Dyson and Mehta (1963); Berry et al. (1977); Berry and Tabor (1976). Short-range statistics, encompassing level statistics, and the ratio of consecutive level spacings reveal the symmetries and randomness in the system Atas et al. (2013). Long-range correlations are captured by the Spectral Form Factor (SFF) Agam et al. (1995); Leviandier et al. (1986); Wilkie and Brumer (1991); Brézin and Hikami (1997), which in quantum chaotic systems, displays a dip-ramp-plateau structure Guhr et al. (1998). It indicates the level-repulsion among the eigenvalues, a characteristic feature of quantum chaos Cotler et al. (2017) and has also been understood from the semiclassical gravity Saad et al. (2018). Recent investigations have yielded precise measurements of SFF, marking a significant advancement in probing quantum chaos in quantum many-body systems through experiments Das et al. (2024); Dong et al. (2024).
Analogously, the upsurge in research has been directed towards unraveling the quantum chaotic properties of open quantum systems, primarily through the Lindbladian dynamics in RMT Denisov et al. (2019); Yang et al. (2024) and operator growth, see Nandy et al. (2024) for a comprehensive review. These systems, interacting with external environments, are often described by non-Hermitian Hamiltonians Bender and Boettcher (1998); Shibata and Katsura (2019); Ashida et al. (2021); Cornelius et al. (2022); Matsoukas-Roubeas et al. (2023a), leading to complex eigenvalues. This connotes a departure from conventional methods suited for Hermitian systems and necessitates new methods for spectral analysis. Advances include generalizing level spacing ratios to complex planes Sá et al. (2020) and adapting SFF to dissipative systems, known as Dissipative Spectral Form Factor (DSFF) Li et al. (2021, 2024); Cipolloni and Grometto (2024), showing a quadratic ramp-plateau pattern diverging from the linear ramp structure seen in Hermitian systems. Several other generalizations for SFF in non-Hermitian and open systems have also been proposed García-García et al. (2023a); Xu et al. (2021); Zhou et al. (2024).
Recently, an alternative approach Kawabata et al. (2023) considers the real and positive singular values of non-Hermitian Hamiltonians instead of the complex eigenvalues, obtained through Singular Value Decomposition (SVD) Nielsen and Chuang (2010), as opposed to diagonalization. Aligning with the Hermitization process Feinberg and Zee (1997), this method facilitates a straightforward generalization of the frameworks used for Hermitian systems, bypassing the complexities associated with the complex plane. Such definitions have been fruitful in defining the singular spacing ratios Kawabata et al. (2023) and singular form factor ($\upsigma$FF) Roccati et al. (2024) in a much elegant manner that effectively captures the integrable to chaotic transitions, non-Hermitian many-body localization Hamazaki et al. (2019); Roccati et al. (2024) and non-Hermitian skin effect Martinez Alvarez et al. (2018); Yao and Wang (2018); Hamanaka and Kawabata (2024). Outside this domain, SVD has also found applications in generalizing entanglement entropy for non-Hermitian transition matrices in AdS/CFT correspondence Parzygnat et al. (2023).
@@
In this paper, we utilize the SVD framework to study the quantum chaotic properties and their transitions in the sparse, non-Hermitian version of the Sachdev-Ye-Kitaev (SYK) model Sachdev and Ye (1993); Kitaev (2015). Standing as a quintessential archetype for quantum chaos and possessing a dual gravitational theory at low temperatures Maldacena and Stanford (2016), the SYK model (see See Chowdhury et al. (2022); Rosenhaus (2019) for comprehensive reviews) can be adapted into sparse variants Xu et al. (2020); Tezuka et al. (2023); Anegawa et al. (2023); Caceres et al. (2021); Cáceres et al. (2023), maintaining its chaotic nature up to a critical sparsity threshold García-García et al. (2021); Orman et al. (2024). This adaptation holds theoretical considerations in holography as well as the practical significance for the deployment of the SYK model on quantum processors Jafferis et al. (2022); Kobrin et al. (2023), where managing a sparse matrix is more viable. Our focus is particularly drawn to the influence that sparsity exerts on the average singular value spacing ratio, denoted by $\langle r_{\upsigma}\rangle$, and the $\upsigma$FF, which serves as an extension of the conventional $r$-value Atas et al. (2013) and the SFF in Hermitian systems, respectively. Additionally, we introduce singular complexity, inspired by spectral complexity, a boundary dual to the volume of the Einstein-Rosen bridge in gravitational theories Iliesiu et al. (2022). Our findings suggest that these metrics proficiently encapsulate the transition from chaos to integrability in the non-Hermitian SYK model, as we outline in the following sections.
@@
non-Hermitian SYK model: The prototypical model we consider is the $4$-body non-Hermitian SYK model (nSYK), given by the following Hamiltonian García-García and Godet (2021); García-García et al. (2022a)
$H_{\mathrm{nSYK}}=\sum_{1\leq a<b<c<d\leq N}(J_{abcd}+iM_{abcd})\,\psi_{a}\psi_{b}\psi_{c}\psi_{d}\,,$ (1)
where $J_{abcd}$ and $M_{abcd}$ are the independent random couplings drawn from the Gaussian ensemble with zero mean and variance $\braket{J_{abcd}^{2}}=\braket{M_{abcd}^{2}}=6/N^{3}$. The variables $\psi_{k}$ are the Majorana fermions obeying Clifford algebra $\{\psi_{a},\psi_{b}\}=\delta_{ab}$. For numerical purposes, we consider the $4$-body interactions only. The couplings $M_{abcd}$ explicitly break the Hermiticity of the Hamiltonian (1), which otherwise reduces to the usual Hamiltonian Sachdev and Ye (1993); Kitaev (2015) when $M_{abcd}$ vanishes for all $\{a,b,c,d\}$. This model has been pivotal in exploring non-Hermitian quantum chaos García-García et al. (2022b); Cipolloni and Kudler-Flam (2023).
In the following, we also consider the sparse version of the Hamiltonian (1), given by
@@
In this sparse formulation, $x_{ijkl}\in[0,1]$ are random variables that introduce an element of sparsity to the Hamiltonian. They follow a Gaussian distribution with zero mean and a variance of $\braket{x^{2}_{abcd}}=6/(pN^{3})$ Xu et al. (2020); García-García et al. (2021). The parameter $p$ governs the probability of $x_{abcd}$ being unity, thus determining the level of sparsity within the Hamiltonian. A fully dense model is characterized by $x_{abcd}=1$ or $p=1$ for all combinations of $\{a,b,c,d\}$, aligning with the original nSYK Hamiltonian (1). Conversely, a completely sparse model would have $x_{abcd}=0$ or $p=0$ for all combinations, representing the other extreme Xu et al. (2020).
The probabilistic nature of $x_{abcd}$ implies that the exact number of non-zero terms in $H_{\mathrm{nSYK}}^{\mathrm{sparse}}$ is not fixed but rather determined by the probability $p$. It is worth noting that one can choose to fix the number of non-zero terms in the sparse Hamiltonian (2) for each realization. However, such distinctions become less significant when considering the average over a large number of ensembles García-García et al. (2021); Orman et al. (2024). For our study, we follow Orman et al. (2024) without fixing the total number of terms in the sparse Hamiltonian in each realization.
@@
Singular value decomposition and singular value statistics: In Hermitian systems, the eigenspectrum of the Hamiltonian is a key indicator of chaotic dynamics. The eigenvalues $E_{n}$ are real, allowing for the definition of consecutive level spacings $s_{n}=E_{n+1}-E_{n}$, as an ordered list. For large Hilbert spaces, the average consecutive level spacing ratio $r_{n}$ is defined as $r_{n}=\mathrm{min}(s_{n},s_{n+1})/\mathrm{max}(s_{n},s_{n+1})$ with $\braket{r}=\mathrm{mean}(r_{n})$ being the average $r$ value Atas et al. (2013). Its usefulness lies in the simplification of the numerical analysis by bypassing the need to unfold the spectrum for level statistics. Typically, the $r$-ratio statistics is derived from the mid-spectrum, avoiding edge anomalies. For Gaussian Orthogonal Ensemble (GOE), Gaussian Unitary Ensemble (GUE), and Gaussian Symplectic Ensemble (GSE), the level statistics follow the Wigner-Dyson distribution with an average $\braket{r}_{\mathrm{GOE}}\approx 0.53590$, $\braket{r}_{\mathrm{GUE}}\approx 0.60266$ and $\braket{r}_{\mathrm{GSE}}\approx 0.67617$ respectively while the Poissonian distribution corresponds to $\braket{r}_{\mathrm{P}}\approx 0.386$ Atas et al. (2013). It is important to note that these values (except for $\braket{r}_{\mathrm{P}})$ are for three-dimensional matrices, and they decrease as the dimension increases. Numerically such values can obtained by fitting the data at appropriate matrix dimensions Atas et al. (2013).
While $\braket{r}$-value is well-defined for Hermitian systems, it falls short in applying for non-Hermitian systems where eigenvalues are complex, and thus an ordered sequence of level spacings is not feasible. In such scenarios, one must consider complex spacing ratios Sá et al. (2020), which are defined in the two-dimensional complex plane. Consequently, the consecutive level spacings are generalized to the Euclidean distance between the eigenvalues in the complex plane, and the complex spacing ratio is then defined by taking the ratio of the distance from a given eigenvalue to its nearest neighbor, to the distance to its second nearest neighbor. However, this definition does not reduce to $\braket{r}$-value for Hermitian systems. Rather it reduces to a different $\braket{r}$-value formed by the closest and second closest neighbors from the ordered set Srivastava et al. (2018).
In a recent work Kawabata et al. (2023), an alternate proposal was suggested. Instead of focusing on complex eigenvalues, the analysis is based on singular values derived from the singular value decomposition (SVD) of the Hamiltonian. SVD is expressed as $H=U\Sigma V^{\dagger}$, where $U$ and $V$ are $N\times N$ unitary matrices satisfying $U^{\dagger}U=V^{\dagger}V=1$, and $\Sigma=\mathrm{diag}(\sigma_{1},\cdots,\sigma_{N})$ is a diagonal matrix containing the singular values $\sigma_{i}$. Singular values are always positive and applicable to both Hermitian and non-Hermitian matrices. For Hermitian matrices, singular values reduce to the absolute values of the eigenvalues, but for non-Hermitian matrices, the singular values and eigenvalues are not straightforwardly related Allard and Kieburg (2024). Importantly, the techniques of SVD are advantageous over other methods such as bi-orthogonalization, where the Hamiltonian can be diagonalized by a bi-orthogonal transformation such that $H=P\Lambda Q^{\dagger}$, where $\Lambda=\mathrm{diag}(E_{1},\cdots E_{n})$ is a diagonal matrix containing the complex eigenvalues of the non-Hermitian Hamiltonian Brody (2013). The matrices $P$ and $Q$ are not individually unitary, but they are bi-unitary satisfying $P^{\dagger}Q=Q^{\dagger}P=1$.
@@
SVD is a fundamental matrix factorization technique in linear algebra Nielsen and Chuang (2010). Curiously, the singular values of a non-Hermitian matrix $H$ can be computed without directly performing SVD. This method, known as Hermitization Feinberg and Zee (1997), involves constructing a new matrix $\mathsf{H}$, twice the dimension of $H$. This larger matrix incorporates both $H$ and its Hermitian conjugate $H^{\dagger}$ in a block structure. The matrix $\mathsf{H}$ is defined as Feinberg and Zee (1997)
@@
The eigenvalues of $\mathsf{H}$ are {±$\{\pm\sigma_{n}\}$, precisely containing the singular values of $H$. This is because the non-zero blocks in $\mathsf{H}$ are off-diagonal, leading to a characteristic polynomial that yields the squares of the singular values of $H$ as eigenvalues. This is equivalent to constructing an effective Hamiltonian $H_{\mathrm{eff}}$, by taking the square root of the product $H$ and $H^{\dagger}$ such that $H_{\mathrm{eff}}=\sqrt{H^{\dagger}H}$ or $H_{\mathrm{eff}}=\sqrt{HH^{\dagger}}$. The eigenvalues of $H_{\mathrm{eff}}$ represent the singular values of $H$ Feinberg and Zee (1997). Such construction of a Hermitized matrix resembles the formulation of the Wishart SYK model and its supersymmetric version, comprised of non-Hermitian conserved charges or supercharges Iyoda et al. (2018); Sá and García-García (2022).
Given the singular values $\{\sigma_{i}\}$ of the corresponding non-Hermitian Hamiltonian $H$, one analogously defines $\lambda_{i}=\sigma_{i+1}-\sigma_{i}$ as the singular value spacings. Consequently, the singular-value-spacing ratio $r_{\upsigma}$ is defined analogous to the $r$-ratio as Kawabata et al. (2023)
$r_{\upsigma,n}=\frac{\mathrm{min}(\lambda_{i},\lambda_{i+1})}{\mathrm{max}(\lambda_{i},\lambda_{i+1})}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \braket{r_{\upsigma}}=\mathrm{mean}(r_{\upsigma,n})\,.$ (4)
Referred to as the average $r_{\upsigma}$ value, these values adhere to statistical distributions akin to those observed for non-Hermitian Hamiltonians, similar to the $\braket{r}$-value distributions for Hermitian matrices. This framework has been instrumental in the $38$-fold symmetry classification in non-Hermitian systems Kawabata et al. (2023).
@@
Table 1: The Table shows the $\braket{r_{\upsigma}}$-values for the non-Hermitian SYK model (1) for different system sizes $N=20\,(10000)$, $N=22\,(6000)$, $N=24\,(2000)$, $N=26\,(1000)$, $N=28\,(100)$, and $N=30\,(50)$, where the parenthesis includes the number of samples taken. Here $N$ mod $8$ = 0 (GOE), 2, 6 (GUE), and 4 (GSE) You et al. (2017). The values are in precise agreement (with numerical accuracy) with the $\braket{r}_{\mathrm{RMT}}$-values for large $N$ results given in Atas et al. (2013).
@@
Table  1 showcases the $\braket{r_{\upsigma}}$-value for the dense non-Hermitian SYK model (1) for varying system sizes $N$, representing the total number of fermions. The data is juxtaposed with corresponding Hermitian Gaussian ensembles, revealing a striking correlation that underscores the utility of the $\braket{r_{\upsigma}}$-value in understanding non-Hermitian systems. Figure 1 correspondingly shows the consecutive singular level-spacing distribution $p_{\sigma}(\lambda)$ for $N=24$ (GOE), $N=26$ (GUE) and $N=28$ (GSE) respectively, which are approximated by the Wigner-Dyson distribution Mehta (1991)
$p(\lambda)=\begin{cases}\frac{\pi\lambda}{2}\,e^{-\frac{\pi\lambda^{2}}{4}}&\leavevmode\nobreak\ \leavevmode\nobreak\ \mathrm{GOE}\,,\\ \frac{32\lambda^{2}}{\pi^{2}}\,e^{-\frac{4\lambda^{2}}{\pi}}&\leavevmode\nobreak\ \leavevmode\nobreak\ \mathrm{GUE}\,,\\ \left(\frac{64}{9\pi}\right)^{3}\,e^{-\frac{64\lambda^{2}}{9\pi}}&\leavevmode\nobreak\ \leavevmode\nobreak\ \mathrm{GSE}\,.\end{cases}$ (5)
[…]
@@
Figure 1: The consecutive singular level-spacing distribution in the non-Hermitian dense SYK model (1) for $N=24\,(2000)$ (GOE), $26\,(1000)$ (GUE) and $28\,(100)$ (GSE) with the corresponding realizations shown in the brackets. The solid lines represent distribution for GOE, GUE, and GSE respectively taking matrices of size $10^{4}\times 10^{4}$ with averaging over 1000 realizations.
SVD in sparse non-Hermitian SYK model: In the preceding discussion, we delved into the behavior of the $r_{\upsigma}$-ratio for the non-Hermitian SYK model, given by equation (1), across various system sizes. We now turn our attention to the sparse variant of this model, which is governed by the sparsity parameter $p$. Our aim is to elucidate the impact of sparsity on the $\braket{r_{\upsigma}}$-ratio during the transition from a densely connected to a sparsely connected regime.
@@
Figure 2 illustrates how the average $\braket{r_{\upsigma}}$-value varies with the sparsity parameter $p$ for different system sizes. Each system size, representing a distinct ensemble, is depicted in a unique color. The dotted lines mark the expected values for Poisson statistics and the Hermitian ensemble, as specified in Table 1. It is observed that the Hamiltonian (2) retains its chaotic nature up to a critical level of sparsity denoted by $p_{\mathrm{crit}}$. This chaotic behavior is also corroborated by the $\braket{r}$-ratio presented in Table 1. Consequently, we define a transition point $p_{\mathrm{crit}}$ beyond which the Hamiltonian (2) ceases to exhibit chaotic dynamics. Such transition point is obtained by considering approximately $99\%$ of the corresponding $\braket{r_{\upsigma}}$ value at $p=1$ Orman et al. (2024). Notably, as the system size increases, this transition point migrates towards lower sparsity levels. When the sparsity is sufficiently low, the $\braket{r_{\upsigma}}$-value aligns with the Poisson limit, signifying the Hamiltonian’s transition to integrability.
@@
In conjunction with the above, we also present the relationship between the critical sparsity $p_{\mathrm{crit}}$ and the system size $N$. Intriguingly, the critical value diminishes as the system size increases. This reduction follows a power-law trend, best described by fitting the data to the curve Orman et al. (2024)
$p=kN\binom{N}{4}^{-1}\approx\frac{24k}{N^{3}}\,,$ (6)

### v2

@@ Abstract
Exploring the spectral properties of non-Hermitian systems presents a substantial theoretical challenge due to the presence of a complex eigenvalue spectrum. Singular values for such systems are inherently real and non-negative, and the techniques used for Hermitian systems can be used with ease. As a prototypical example of such systems, we investigate the singular-value spectrum of a non-Hermitian extension of the sparse Sachdev-Ye-Kitaev (SYK) model, a solvable toy model of quantum chaos and quantum gravity with significant interest in digital quantum simulation. Our findings reveal a congruence between the statistics of singular values and those of the analogous Hermitian Gaussian ensembles. An increase in sparsity results in the model deviating from its chaotic behavior, a phenomenon precisely captured by both short and long-range correlations in the singular-value spectrum. These findings indicate the existence of a critical sparsity threshold, beyond which the chaotic nature of the model breaks down, suggesting a potential collapse of holographic duality in such systems.
@@
Introduction: Spectral statistics serve as a critical tool for probing the energy levels in quantum systems, offering insights into the dynamics within quantum chaos and Random Matrix Theory (RMT) Mehta (1991); Bohigas et al. (1984); Wigner (1955, 1957); Dyson (1962a, b); Berry et al. (1977); Berry and Tabor (1976). Short-range level statistics such as the ratio of consecutive level spacings Atas et al. (2013) and the long-range correlations such as the Spectral Form Factor (SFF) Agam et al. (1995); Guhr et al. (1998); Leviandier et al. (1986); Wilkie and Brumer (1991); Brézin and Hikami (1997); Cotler et al. (2017) reveal the symmetries and randomness in the system. The latter has also been understood from the semiclassical gravity Saad et al. (2018) and measured in many-body scenarios Das et al. (2024); Dong et al. (2024). Analogously, the upsurge in research has been directed towards unraveling the quantum chaotic properties of open quantum systems Denisov et al. (2019); Yang et al. (2024); Nandy et al. (2024) described by non-Hermitian Hamiltonians Bender and Boettcher (1998); Shibata and Katsura (2019); Ashida et al. (2021); Cornelius et al. (2022); Matsoukas-Roubeas et al. (2023a). This connotes a departure from conventional methods suited for Hermitian systems and necessitates new methods for spectral analysis Sá et al. (2020); Li et al. (2021, 2024); Cipolloni and Grometto (2024); García-García et al. (2023a); Xu et al. (2021); Zhou et al. (2024). One such method approaches through the Singular Value Decomposition (SVD) Kawabata et al. (2023), which facilitates a straightforward generalization of the frameworks used for Hermitian systems with real and non-negative singular values. Consequently, this enables us to define the singular spacing ratios Kawabata et al. (2023) and singular form factor ($\upsigma$FF) Roccati et al. (2024) in a much more elegant manner capturing transitions from integrability to chaos, many-body localization Hamazaki et al. (2019); Roccati et al. (2024), and the non-Hermitian skin effect Martinez Alvarez et al. (2018); Yao and Wang (2018); Hamanaka and Kawabata (2024). Outside this domain, SVD has also found applications in generalizing entanglement entropy for non-Hermitian transition matrices in AdS/CFT correspondence Parzygnat et al. (2023).
@@
In this paper, we utilize the SVD framework to study the quantum chaotic properties and their transitions in the sparse, non-Hermitian version of the Sachdev-Ye-Kitaev (SYK) model Sachdev and Ye (1993); Kitaev (2015). Standing as a quintessential archetype for quantum chaos and possessing a dual gravitational theory at low temperatures Maldacena and Stanford (2016), the SYK model (see See Chowdhury et al. (2022); Rosenhaus (2019) for comprehensive reviews) can be adapted into sparse variants Xu et al. (2020); Tezuka et al. (2023); Anegawa et al. (2023); Caceres et al. (2021); Cáceres et al. (2023), maintaining its chaotic nature up to a critical sparsity threshold García-García et al. (2021); Orman et al. (2024). This adaptation holds theoretical considerations in holography as well as the practical significance for the deployment of the SYK model on quantum processors Jafferis et al. (2022); Kobrin et al. (2023), where managing a sparse matrix is more viable. Our focus is particularly drawn to the influence that sparsity exerts on the average singular value spacing ratio, denoted by $\langle r_{\upsigma}\rangle$, and the $\upsigma$FF, which serves as an extension of the conventional $\braket{r}$-value Atas et al. (2013) and the SFF in Hermitian systems, respectively. Additionally, we introduce singular complexity, inspired by spectral complexity, a boundary dual to the volume of the Einstein-Rosen bridge in gravitational theories Iliesiu et al. (2022). Our findings suggest that these metrics proficiently encapsulate the transition from chaos to integrability in the non-Hermitian SYK model, as we outline in the following sections.
@@
Non-Hermitian sparse SYK model: The prototypical model we consider is the $4$-body non-Hermitian sparse SYK model (nSYK) Sachdev and Ye (1993); Kitaev (2015), given by the following Hamiltonian
@@
where $J_{abcd}$ and $M_{abcd}$ are the independent random couplings drawn from the Gaussian ensemble with zero mean and variance $\braket{J_{abcd}^{2}}=\braket{M_{abcd}^{2}}=6/N^{3}$ (see Lau et al. (2021); Ozaki and Katsura (2025) for clean SYK model). The couplings $M_{abcd}$ explicitly break the Hermiticity of the Hamiltonian. The variables $\psi_{k}$ are the Majorana fermions obeying Clifford algebra $\{\psi_{a},\psi_{b}\}=\delta_{ab}$. Additionally, the random variables $x_{ijkl}\in[0,1]$ introduce an element of sparsity to the Hamiltonian. They follow a Gaussian distribution with zero mean and a variance of $\braket{x^{2}_{abcd}}=6/(pN^{3})$ Xu et al. (2020); García-García et al. (2021). The parameter $p$ governs the probability of $x_{abcd}$ being unity, thus determining the sparsity level within the Hamiltonian. A fully dense model is characterized by $x_{abcd}=1$ or $p=1$ for all combinations of $\{a,b,c,d\}$, aligning with the original nSYK Hamiltonian García-García and Godet (2021); García-García et al. (2022a, b); Cipolloni and Kudler-Flam (2023). Conversely, a completely sparse model would have $x_{abcd}=0$ or $p=0$ for all combinations, representing the other extreme Xu et al. (2020). The addition of sparsity effectively reduces the total number of independent terms in the Hamiltonian from $O(N^{4})$ to $O(pN^{4})\sim O(kN)$, where $k$ is defined in (4). Yet, the probabilistic nature of $x_{abcd}$ implies that the exact number of non-zero terms in $H_{\mathrm{nSYK}}^{\mathrm{sparse}}$ is not fixed but rather determined by the probability $p$. We follow Orman et al. (2024) without fixing the total number of terms in the sparse Hamiltonian in each realization.
@@
Singular value decomposition and singular value statistics: In Hermitian systems, the eigenspectrum of the Hamiltonian is a key indicator of chaotic dynamics. The eigenvalues $E_{n}$ are real, allowing for the definition of consecutive level spacings $s_{n}=E_{n+1}-E_{n}$, as an ordered list. For large Hilbert spaces, the average consecutive level spacing ratio $r_{n}$ is defined as $r_{n}=\mathrm{min}(s_{n},s_{n+1})/\mathrm{max}(s_{n},s_{n+1})$ with $\braket{r}=\mathrm{mean}(r_{n})$ being the average $r$ value Atas et al. (2013). Its usefulness lies in simplifying the numerical analysis by bypassing the need to unfold the spectrum for level statistics. While $\braket{r}$-value is well-defined for Hermitian systems, it falls short in applying for non-Hermitian systems where eigenvalues are complex, and thus an ordered sequence of level spacings is not feasible. In such scenarios, one may consider complex spacing ratios Sá et al. (2020) and correspondingly the statistics of complex eigenvalues. However, an alternate proposal suggests focusing on singular values derived from the singular value decomposition (SVD) of the Hamiltonian Kawabata et al. (2023). SVD is expressed as $H=U\Sigma V^{\dagger}$, where $U$ and $V$ are unitary matrices, and $\Sigma$ is a diagonal matrix containing the real and non-negative singular values $\sigma_{i}$ Nielsen and Chuang (2010). In Hermitian matrices, singular values are simply the absolute values of the eigenvalues, whereas, in non-Hermitian matrices, their relationship is more complex Allard and Kieburg (2024).
@@
The singular values of a non-Hermitian matrix $H$ can be computed by the Hermitization process Feinberg and Zee (1997), which involves constructing a new matrix $\mathsf{H}$, twice the dimension of $H$. This larger matrix incorporates both $H$ and its Hermitian conjugate $H^{\dagger}$ in a block structure Feinberg and Zee (1997)
@@
The eigenvalues of $\mathsf{H}$ are {±$\{\pm\sigma_{n}\}$, precisely containing the singular values of $H$. This is equivalent to constructing an effective Hamiltonian $H_{\mathrm{eff}}=\sqrt{H^{\dagger}H}$ or $H_{\mathrm{eff}}=\sqrt{HH^{\dagger}}$ so that the eigenvalues of $H_{\mathrm{eff}}$ represent the singular values of $H$ Feinberg and Zee (1997). Such construction of a Hermitized matrix resembles the formulation of the Wishart SYK model and its supersymmetric version, comprised of non-Hermitian conserved charges or supercharges Li et al. (2017); Sá and García-García (2022).
Given the singular values $\{\sigma_{i}\}$ of the corresponding non-Hermitian Hamiltonian $H$, one analogously defines $\lambda_{n}=\sigma_{n+1}-\sigma_{n}$ as the singular value spacings. Consequently, the singular-value-spacing ratio $r_{\upsigma,n}$ is defined analogous to the $r$-ratio as Kawabata et al. (2023)
$r_{\upsigma,n}=\frac{\mathrm{min}(\lambda_{n},\lambda_{n+1})}{\mathrm{max}(\lambda_{n},\lambda_{n+1})}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \braket{r_{\upsigma}}=\mathrm{mean}(r_{\upsigma,n})\,.$ (3)
Referred to as the $\braket{r_{\upsigma}}$-value, they adhere to statistical distributions akin to those observed for non-Hermitian Hamiltonians, similar to the $\braket{r}$-value distributions for Hermitian matrices. This framework has been instrumental in addressing the $38$-fold symmetry classification in non-Hermitian systems Kawabata et al. (2023).
@@
Table 1: The $\braket{r_{\upsigma}}$-values for the non-Hermitian dense SYK model for different system sizes $N=20\,(10000)$, $N=22\,(6000)$, $N=24\,(2000)$, $N=26\,(1000)$, $N=28\,(100)$, and $N=30\,(50)$, where the parenthesis includes the number of samples taken. Based on the $\braket{r_{\upsigma}}$ values, we identify the periodicity as $N$ mod 8 = 0 (GOE), 2, 6 (GUE), and 4 (GSE) You et al. (2017). These GOE, GUE, and GSE correspond to the eigenvalue statistics of the Hermitian matrices.
The values are in precise agreement (within numerical accuracy) with the $\braket{r}_{\mathrm{RMT}}$-values for large $N$ results in Atas et al. (2013).
@@
Table  1 showcases the $\braket{r_{\upsigma}}$-value for the dense non-Hermitian SYK model for varying system sizes $N$, representing the total number of fermions. The data is juxtaposed with corresponding eigenvalue statistics of Hermitian Gaussian ensembles, revealing a striking correlation emphasizing the utility of the $\braket{r_{\upsigma}}$-value in non-Hermitian systems. Similar to its Hermitian counterpart, we observe an $N$ mod 8 periodicity using singular value statistics: $N$ mod 8 = 0 corresponds to GOE, $N$ mod 8 = 2 and 6 corresponds to GUE, and $N$ mod 8 = 4 corresponds to GSE for the eigenvalue statistics of the Hermitian matrices. Figure 1 correspondingly shows the consecutive singular level-spacing distribution $p_{\sigma}(\lambda)$ for $N=24$ (GOE), $N=26$ (GUE) and $N=28$ (GSE) respectively, which are approximated by the analogous Hermitian Wigner-Dyson distribution Mehta (1991). Note that we are comparing the singular value statistics of non-Hermitian matrices to the eigenvalue statistics of Hermitian matrices. The singular value correlations in Hermitian matrices are generally weaker than the eigenvalue correlations and exhibit different spacing ratios Kawabata et al. (2023). However, the nSYK falls into the different non-Hermitian symmetry classes, which can also be used to show the validity of the singular value statistics García-García et al. (2022a).
@@
Figure 1: The consecutive singular level-spacing distribution in the non-Hermitian dense SYK model for $N=24\,(2000)$ (GOE), $26\,(1000)$ (GUE), and $28\,(100)$ (GSE) with the corresponding realizations shown in the brackets. The solid lines represent distribution for GOE, GUE, and GSE respectively taking matrices of size $10^{4}\times 10^{4}$ with averaging over 1000 realizations.
@@
SVD in sparse non-Hermitian SYK model: We now turn our attention to the sparse variant of the model, which is governed by the sparsity parameter $p$. Figure 2 illustrates the dependence of $p$ on $\braket{r_{\upsigma}}$-value for different system sizes. Each system size, representing a distinct ensemble, is depicted in a unique color. The dotted lines mark the expected values for Poisson statistics and the Hermitian ensemble, as specified in Table 1. It is observed that the Hamiltonian (1) retains its chaotic nature up to a critical level of sparsity denoted by $p_{\mathrm{crit}}$. This chaotic behavior is also corroborated by the $\braket{r_{\upsigma}}$-ratio presented in Table 1. Consequently, we define a transition point $p_{\mathrm{crit}}$ beyond which the Hamiltonian (1) ceases to exhibit chaotic dynamics. Such transition point is obtained by considering approximately $99\%$ of the corresponding $\braket{r_{\upsigma}}$ value at $p=1$ Orman et al. (2024). Notably, as the system size increases, this transition point migrates towards lower sparsity levels. When the sparsity is sufficiently low, the $\braket{r_{\upsigma}}$-value aligns with the Poisson limit, signifying the Hamiltonian’s transition to complete integrability.Note that when a term of the Hamiltonian very weakly breaks a degeneracy, the resulting spectrum has $r_{\sigma}$ values close to zero, which can cause the $\braket{r_{\upsigma}}$ value to fall below the Poisson limit.
@@
In conjunction with the above, the relationship between the critical sparsity $p_{\mathrm{crit}}$ and the system size $N$ is also noteworthy. The critical value diminishes as the system size increases. This reduction follows a power-law trend, best described by fitting the data to the curve Orman et al. (2024)
$p_{\mathrm{crit}}=kN\binom{N}{4}^{-1}\approx\frac{24k}{N^{3}}\,,$ (4)